# Competing effect of disorder on phase separation in active systems: from facilitated to hindered growth and rough interfaces

### v1

@@ A.3 Fractal dimension of interface
To investigate the fractal nature of the domains we plot the fractal dimension of the interface $d_{f}$ vs. disorder $\rho_{d}$ in Fig.8 (c) (main manuscript). To calculate the fractal dimension of the interface we follows the below procedure: For a fractal interface, the length of the interface decreases with time $t$ as $B(t)\approx t^{-\frac{d_{f}-1}{3}}$, where, $d_{f}$ is the fractal dimension Streitenberger (2000). The length of the interface is determined using a box-counting algorithm. We begin by evaluating the fluctuation of the order parameter at each lattice point $(i,j)$ as $\delta\rho(i,j)=\rho(i,j)-\rho_{0}$ where, $\rho_{0}$ is the mean value of $\rho$. A point is classified as part of a high-density region if $\delta\rho>0$, otherwise, it is considered to belong to a low-density region. Each lattice point is then further categorized based on the number of its high-density neighbors, denoted by by $n_{h}$. Points with $n_{h}=4$ are considered core points, residing within clusters. Points with $2\leq n_{h}\leq 3$ that has at least one neighbor which is a core point are classified as edge points, lying at the interface. Points outside these criteria are deemed isolated, located in low-density regions. In this study, the spatial grid size is uniform in both $x$ and $y$-directions. The perimeter or length of the interface, $B(t)$, is then calculated by multiplying the number of edge points by the spatial grid size.
A.4 Density fluctuation
@@
To calculate the density fluctuations in the system $\Delta\rho_{n}$, we divide the simulation box of size $512\times 512$ into square blocks of different sizes from $1\times 1$ to $256\times 256$. For each block size, we calculate the total density within each block across multiple time steps in the steady state and determine the average density $\rho_{n}$, so as the fluctuates from that average. By repeating this process for different block sizes, we calculate average densities $\rho_{n}$ and their variances $\Delta\rho_{n}$ for different system with $\rho_{d}=0-0.5$

### v2

@@ Appendix E Fractal dimension of interface
To investigate the fractal nature of the domains we plot the fractal dimension of the interface $d_{f}$ vs. disorder $\rho_{d}$ in Fig.7 (c) (main manuscript). To calculate the fractal dimension of the interface we follows the below procedure: For a fractal interface, the length of the interface decreases with time $t$ as $B(t)\approx t^{-\frac{d_{f}-1}{3}}$, where, $d_{f}$ is the fractal dimension [63]. The length of the interface is determined using a box-counting algorithm. We begin by evaluating the fluctuation of the order parameter at each lattice point $(i,j)$ as $\delta\rho(i,j)=\rho(i,j)-\rho_{0}$ where, $\rho_{0}$ is the mean value of $\rho$. A point is classified as part of a high-density region if $\delta\rho>0$, otherwise, it is considered to belong to a low-density region. Each lattice point is then further categorized based on the number of its high-density neighbors, denoted by by $n_{h}$. Points with $n_{h}=4$ are considered core points, residing within clusters. Points with $2\leq n_{h}\leq 3$ that has at least one neighbor which is a core point are classified as edge points, lying at the interface. Points outside these criteria are deemed isolated, located in low-density regions. In this study, the spatial grid size is uniform in both $x$ and $y$-directions. The perimeter or length of the interface, $B(t)$, is then calculated by multiplying the number of edge points by the spatial grid size.
Appendix F Density fluctuation
@@
To calculate the density fluctuations in the system $\Delta\rho_{n}=\sqrt{\langle\rho^{2}\rangle-\langle\rho\rangle^{2}}$, we divide the simulation box of size $512\times 512$ into square blocks of different sizes from $1\times 1$ to $256\times 256$. For each block size, we calculate the total density within each block across multiple time steps in the steady state and determine the average density $\rho_{n}$, so as the fluctuations from that average. By repeating this process for different block sizes, we calculate average densities $\rho_{n}$ and their variances $\Delta\rho_{n}$ for different system with $\rho_{d}=0-0.5$
Appendix G Description of Movies
Movie1(MV1)- In MV1, we show the early time evolution of the local density field $\rho(r)$ of the active particles for system size $K=128$ and disorder density $\rho_{d}=0.05$. The real time is displayed on the top of each snapshots. The black dots represents the pinning sites or disorder sites. The color bar presents the magnitude of the local density field at the lattice points.
Movie2(MV2)- In MV2, we illustrates the late time evolution of the local density field $\rho(r)$ for same system (Real time is mentioned on the top). And the black dots and color bar represents the same as in MV1. The first two prominent clusters are identified using the numbers (1-2) in the movie.
Movie3(MV3)- The MV3, shows the very late time evolution of the local density field $\rho(r)$ for same system (Real time is mentioned on the top).
Movie4(MV4)- The MV4, showcases the time evolution of the local density field $\rho(r)$ for system having $\rho_{d}=0.4$ (Real time is mentioned on the top). The other details are the same as for the above MV1.